OpenStudy (anonymous):

Let a and n be two positive integers with a,n => 2. Prove that if a^n - 1 is prime then a = 2 and n is prime.

OpenStudy (anonymous):

\[a,n \ge 2\]

OpenStudy (anonymous):

what does _> INDICATE

OpenStudy (anonymous):

wrote it above just now

OpenStudy (anonymous):

is equal to or greater than

OpenStudy (anonymous):

yes mate

OpenStudy (anonymous):

it is prime

OpenStudy (anonymous):

i need to prove it

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

any ideas anyone?

OpenStudy (anonymous):

2 power anything - 1 is always prime try it.

OpenStudy (anonymous):

sorry 2 power any prime number.

OpenStudy (anonymous):

but i need to prove it for all numbers aboove and equal to 2

OpenStudy (zarkon):

2^11-1 is not prime

OpenStudy (zarkon):

this is actually really easy...where are you stuck?

OpenStudy (anonymous):

please show how to prove.

OpenStudy (anonymous):

I need to prove for all numbers equal to and above the number 2, that when n is prime and a is 2, the equation \[2^{n}-1\] is always prime. how do i prove this?

OpenStudy (zarkon):

\[2^n-1\] is not always prime...see my counter example above.

OpenStudy (anonymous):

then im lost, dont get the question, my first post shows the whole question as its written

OpenStudy (zarkon):

that one is stated correctly...your recent statement is not

OpenStudy (anonymous):

this question confuses me, can you explain what im supposed to do please?

OpenStudy (zarkon):

is says that if \[2^n-1\] is prime then a must be 2 and n must be prime

OpenStudy (anonymous):

yes i need to prove that

OpenStudy (anonymous):

i get my mistake previously i was working the wrong way around, sorry

OpenStudy (zarkon):

start by showing that is \[2^n-1\] is prime then n is prime

OpenStudy (anonymous):

how do i show that?

OpenStudy (zarkon):

what happens if n is not prime

OpenStudy (anonymous):

when n is not prime the equation is not prime

OpenStudy (anonymous):

but that doesn't prove it for all results?

OpenStudy (zarkon):

prove it is not prime..one step at a time...geesh

OpenStudy (anonymous):

can you explain it fully?

OpenStudy (zarkon):

if n is not prime then n=x*y \[2^n-1=2^{xy}-1=(2^x)^y-1\] this is factorable...you can show that. thus n must be prime

OpenStudy (zarkon):

after you show it is factorable... assume a is odd then a^n-1 is even and not prime so it is not prime so a must be even assume a>2 then a=2*m show \[(2m)^n-1\] is divisible by (2m-1) thus a=2 is the only number that works

OpenStudy (anonymous):

thanks very much i get the proving by contradiction. i dont get how to show the last part, that (2m)^n -1 is divisible by (2m-1). appreciate the help

OpenStudy (zarkon):

prove it by induction

OpenStudy (anonymous):

proving by induction, subbed in n = 1 for base which works, then n+1 so (2m)^(n+1) -1 but dont know how to prove its divisible from there

OpenStudy (zarkon):

or use \[x^n-1=(x-1)(x^{n-1}+x^{n-2}+\cdots+x+1)\]

OpenStudy (anonymous):

id rather use induction, need to learn how to do that for my test

OpenStudy (zarkon):

\[(2m)^{n+1}-1=(2m)^n(2m)-1=(2m)^n(1+(2m-1))-1\] \[=(2m)^n-1+(2m-1)(2m)^n\]

OpenStudy (anonymous):

is that the whole proof?

OpenStudy (zarkon):

you then state that \[(2m-1)|((2m)^n-1)\] by our assumption and \[(2m-1)|(2m-1)(2m)^n\] since 2m-1 is a factor. Thus it dives the sum.

OpenStudy (zarkon):

*divides

OpenStudy (anonymous):

thanks i get it now. how does this show that a = 2 is the only number that works when the equation is prime?